# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

### v1

@@ B.3.3 Video Demos
Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation PolarNet [61] $69.4$ $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$ JS3C-Net [55] $73.6$ $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$ PMF [66] $77.0$ $82.0$ $40.0$ $81.0$ $88.0$ $64.0$ $79.0$ $80.0$ $76.0$ $81.0$ $67.0$ $97.0$ $68.0$ $78.0$ $74.0$ $90.0$ $88.0$ Cylinder3D [65] $77.2$ $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $77.2$ $73.4$ $84.6$ $69.1$ $97.7$ $70.2$ $80.3$ $75.5$ $90.4$ $87.6$ AMVNet [31] $77.3$ $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $84.9$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $91.5$ $88.7$ SPVCNN [47] $77.4$ $80.0$ $30.0$ $91.9$ $90.8$ $64.7$ $79.0$ $75.6$ $70.9$ $81.0$ $74.6$ $97.4$ $69.2$ $80.0$ $76.1$ $89.3$ $87.1$ AF2S3Net [9] $78.3$ $78.9$ $52.2$ $89.9$ $84.2$ $77.4$ $74.3$ $77.3$ $72.0$ $83.9$ $73.8$ $97.1$ $66.5$ $77.5$ $74.0$ $87.7$ $86.8$ 2D3DNet [15] $80.0$ $83.0$ $59.4$ $88.0$ $85.1$ $63.7$ $84.4$ $82.0$ $76.0$ $84.8$ $71.9$ $96.9$ $67.4$ $79.8$ $76.0$ $92.1$ $89.2$ RangeFormer [24] $80.1$ $85.6$ $47.4$ $91.2$ $90.9$ $70.7$ $84.7$ $77.1$ $74.1$ $83.2$ $72.6$ $97.5$ $70.7$ $79.2$ $75.4$ $91.3$ $88.9$ GASN [58] $80.4$ $85.5$ $43.2$ $90.5$ $92.1$ $64.7$ $86.0$ $83.0$ $73.3$ $83.9$ $75.8$ $97.0$ $71.0$ $81.0$ $\mathbf{77.7}$ $91.6$ $\mathbf{90.2}$ 2DPASS [56] $80.8$ $81.7$ $55.3$ $92.0$ $91.8$ $73.3$ $86.5$ $78.5$ $72.5$ $84.7$ $75.5$ $97.6$ $69.1$ $79.9$ $75.5$ $90.2$ $88.0$ LidarMultiNet [57] $81.4$ $80.4$ $48.4$ $94.3$ $90.0$ $71.5$ $87.2$ $\mathbf{85.2}$ $\mathbf{80.4}$ $86.9$ $74.8$ $97.8$ $67.3$ $80.7$ $76.5$ $92.1$ $89.6$ SphereFormer [27] $81.9$ $83.3$ $39.2$ $\mathbf{94.7}$ $\mathbf{92.5}$ $77.5$ $84.2$ $84.4$ $79.1$ $\mathbf{88.4}$ $\mathbf{78.3}$ $\mathbf{97.9}$ $69.0$ $\mathbf{81.5}$ $77.2$ $\mathbf{93.4}$ $\mathbf{90.2}$ Fast-FRNet $82.1$ $85.3$ $62.9$ $92.1$ $91.5$ $76.5$ $87.7$ $75.9$ $73.2$ $86.2$ $76.0$ $97.8$ $\mathbf{71.9}$ $80.9$ $77.1$ $90.8$ $88.1$ FRNet $82.5$ $85.8$ $65.4$ $92.1$ $91.6$ $77.4$ $87.9$ $77.4$ $74.3$ $86.0$ $75.7$ $97.8$ $71.8$ $80.8$ $77.0$ $91.0$ $88.3$ UniSeg [32] $\mathbf{83.5}$ $\mathbf{85.9}$ $\mathbf{71.2}$ $92.1$ $91.6$ $\mathbf{80.5}$ $\mathbf{88.0}$ $80.9$ $76.0$ $86.3$ $76.7$ $97.7$ $71.8$ $80.7$ $76.7$ $91.3$ $88.8$
@@
Table 8: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [49] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
Appendix C Broader Impact
C.1 Positive Societal Impacts
C.2 Potential Limitations
Appendix D Public Resources Used
D.1 Public Codebase Used
D.2 Public Datasets Used
D.3 Public Implementations Used
Table 9: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [37] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@ D.3 Public Implementations Used
Figure 8: Failure case examples in FRNet. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
Figure 9: Qualitative comparisons with state-of-the-art LiDAR segmentation methods [4] on the val set of nuScenes [13]. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.

### v2

@@ VI-A Scalability and Robustness
Method (year) mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation PolarNet [33] [’20] $69.4$ $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$ JS3C-Net [85] [’21] $73.6$ $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$ PMF [91] [’21] $77.0$ $82.0$ $40.0$ $81.0$ $88.0$ $64.0$ $79.0$ $80.0$ $76.0$ $81.0$ $67.0$ $97.0$ $68.0$ $78.0$ $74.0$ $90.0$ $88.0$ Cylinder3D [16] [’21] $77.2$ $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $77.2$ $73.4$ $84.6$ $69.1$ $97.7$ $70.2$ $80.3$ $75.5$ $90.4$ $87.6$ AMVNet [20] [’20] $77.3$ $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $84.9$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $91.5$ $88.7$ SPVCNN [15] [’20] $77.4$ $80.0$ $30.0$ $91.9$ $90.8$ $64.7$ $79.0$ $75.6$ $70.9$ $81.0$ $74.6$ $97.4$ $69.2$ $80.0$ $76.1$ $89.3$ $87.1$ AF2S3Net [88] [’21] $78.3$ $78.9$ $52.2$ $89.9$ $84.2$ $77.4$ $74.3$ $77.3$ $72.0$ $83.9$ $73.8$ $97.1$ $66.5$ $77.5$ $74.0$ $87.7$ $86.8$ 2D3DNet [92] [’21] $80.0$ $83.0$ $59.4$ $88.0$ $85.1$ $63.7$ $84.4$ $82.0$ $76.0$ $84.8$ $71.9$ $96.9$ $67.4$ $79.8$ $76.0$ $92.1$ $89.2$ RangeFormer [32] [’23] $80.1$ $85.6$ $47.4$ $91.2$ $90.9$ $70.7$ $84.7$ $77.1$ $74.1$ $83.2$ $72.6$ $97.5$ $70.7$ $79.2$ $75.4$ $91.3$ $88.9$ GASN [90] [’22] $80.4$ $85.5$ $43.2$ $90.5$ $92.1$ $64.7$ $86.0$ $83.0$ $73.3$ $83.9$ $75.8$ $97.0$ $71.0$ $81.0$ $\mathbf{77.7}$ $91.6$ $\mathbf{90.2}$ 2DPASS [67] [’22] $80.8$ $81.7$ $55.3$ $92.0$ $91.8$ $73.3$ $86.5$ $78.5$ $72.5$ $84.7$ $75.5$ $97.6$ $69.1$ $79.9$ $75.5$ $90.2$ $88.0$ LidarMultiNet [93] [’23] $81.4$ $80.4$ $48.4$ $94.3$ $90.0$ $71.5$ $87.2$ $\mathbf{85.2}$ $\mathbf{80.4}$ $86.9$ $74.8$ $97.8$ $67.3$ $80.7$ $76.5$ $92.1$ $89.6$ SphereFormer [68] [’23] $81.9$ $83.3$ $39.2$ $\mathbf{94.7}$ $\mathbf{92.5}$ $77.5$ $84.2$ $84.4$ $79.1$ $\mathbf{88.4}$ $\mathbf{78.3}$ $\mathbf{97.9}$ $69.0$ $\mathbf{81.5}$ $77.2$ $\mathbf{93.4}$ $\mathbf{90.2}$ Fast-FRNet $82.1$ $85.3$ $62.9$ $92.1$ $91.5$ $76.5$ $87.7$ $75.9$ $73.2$ $86.2$ $76.0$ $97.8$ $\mathbf{71.9}$ $80.9$ $77.1$ $90.8$ $88.1$ FRNet $82.5$ $85.8$ $65.4$ $92.1$ $91.6$ $77.4$ $87.9$ $77.4$ $74.3$ $86.0$ $75.7$ $97.8$ $71.8$ $80.8$ $77.0$ $91.0$ $88.3$ UniSeg [23] [’23] $\mathbf{83.5}$ $\mathbf{85.9}$ $\mathbf{71.2}$ $92.1$ $91.6$ $\mathbf{80.5}$ $\mathbf{88.0}$ $80.9$ $76.0$ $86.3$ $76.7$ $97.7$ $71.8$ $80.7$ $76.7$ $91.3$ $88.8$
@@
TABLE IX: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [63] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE X: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [64] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
VI-B Per-Class Results
VI-B1 SemanticKITTI
Tab. VI shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [28]. We compare FRNet with different LiDAR representations, including point-view, range-view, sparse-voxel-view, and multi-view methods. Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods. Compared with SphereFormer [68] and UniSeg [23], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy. More results about efficiency and accuracy can be found in the main body. It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
VI-B2 nuScenes
Tab. VII and Tab. VIII show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [43], respectively. The results demonstrate the great advantages of Fast-FRNet and FRNet. For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters. Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
VI-B3 ScribbleKITTI
To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [63], including RangeNet++ [30], SalsaNext [57], FIRNet [55], CENet [58], and RangeViT [31]. All the models are trained with their official settings. As shown in Tab. IX, FRNet achieves state-of-the-art performance. However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
VI-B4 SemanticPOSS
We also reimplement some popular sparse-voxel-view methods, including SPVCNN [15], Cylinder3D [16], and MinkNet [17], on SemanticPOSS [64]. All the experiments are conducted based on MMDetection3D framework [72]. Tab. X summarizes the class-wise IoU scores among these methods and some other popular range-view methods. Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods. For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.
VII Broader Impact
VII-A Positive Societal Impacts
VII-B Potential Limitations
VIII Public Resources Used
VIII-A Public Codebase Used
VIII-B Public Datasets Used
VIII-C Public Implementations Used